OpenStudy (anonymous):

simplify. (2x^3y^2-x^3y^3) + (6x^2y^3+7-5x^3y^3)

OpenStudy (anonymous):

Match the common terms \[(2x^3y^2-x^3y^3) + (6x^2y^3+7-5x^3y^3) = \] \[(2x^3y^2 + 6x^2y^3+7-5x^3y^3 - x^3y^3) = \] \[(2x^3y^2 + 6x^2y^3-6x^3y^3 + 7) = \] \[[(x^2y^2)(2x + 6y -6xy)] + 7 \]

OpenStudy (anonymous):

7+2y^2x^3+6x^2y^3-6x^3y^3

OpenStudy (anonymous):

the person has already put an explanation so, I jst wrote the answer.

OpenStudy (anonymous):

thank you guys!